IF ONE SIDE IS 66 AND THE OTHER IS 80 WHAT IS THE MEASUREMENT OF THE THIRD SIDE?

Answer by 7asho0m(2)


180 - (side 1 + side 2) = side 3
= 180 - (66 + 80)
= 180 - 146 = 34
Side 3 = 34 degrees